A Bibliography of Supercomputing ’94

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 October 2010 Version 1.03

Abstract

This bibliography records articles presented at the Supercomputing ’94 conference.

Title word cross-reference

N [APG94].

-Body [GKS94, APG94].

1 [AGZ94]. 1-D [AGZ94]. 15 [SS94]. 1994 [IEE94].

3D [JWT^{+}94].

4 [SSSK94].

65 [SHM^{+}94].

9076 [Mra94]. 90D [PHHF94]. ’94 [IEE94].

Abstract [SW94]. Access [KN94]. Achievement [MFI^{+}94]. Acyclic [PM94]. AD [SS94]. Adaptive [SPMH94]. Advanced [Fah94]. Affinity [SE94]. Agenda [FR94]. Algorithm [AGZ94]. Algorithms [DGKL94]. Allocation [LLWN94]. Analysis [BTS94]. and/or [TNC94]. APL [CK94]. Application [PCB94]. Application-Specific [FLRS94]. Applications [BY94]. Applied [JP94]. Applying [Ben94]. Approach [PCB94]. Architecture [UIT94]. Architectures [GMB94]. Array [SW94]. Arrays [NHL94]. Astrophysical [MTES94]. ATM [LHDT94]. ATM-Based [HKM94].

Barnes [GKS94]. Based [PCB94, HKM94]. Basis [SOGD94]. Benchmarks [Gin94].

Benefits [McR94]. Best [Gin94]. Beyond [McR94]. Binding [GC94]. Biomedical [Fah94]. Block [DGKL94]. Block-Recursive [DGKL94]. body [APG94]. Bound [Eck94]. Boundary [WGW^{+}94]. Branch [Eck94]. Buffers [POVDF94]. Building [BGPS94].

Cache [KSF^{+}94]. Case [LHDM94]. Cedar [TV94]. Centre [WMTB94]. CFD [BY94]. Channel [LHDM94]. Chant [HCM94]. Characteristics [KN94]. characterization [PP94]. Chemically [SHM^{+}94]. Cholesky [RS94]. Cluster [CCFH94]. Coherence [KMB94]. Collection [KMY94]. Collective [KBHA94]. Communication [BGPS94]. Communications [KBHA94, Lin94]. Compacting [POVDF94]. Compile [SPMH94]. Compile-time [SPMH94]. Compiler [CK94]. Compiler-Directed [CY94]. Computational [MM94]. Computations [BJTW94]. Computer [FR94]. Computers [KMY94]. Computing [Ben94]. Conjugate [TO94]. Connection [CWM94]. contiguous [LLWN94]. Control [Eck94]. Controlling [WCL94]. Cross [ZI94]. Cross-Loop [ZI94]. Cycles [CL94].

D [AGZ94]. DAMQ [POVDF94]. Data [TT94]. DC [IEE94]. Debugger [CH94]. Dependence [ZI94]. Dependencies [ZI94]. Design [HKM94]. Designing [DGKL94]. Determination [JWT^{+}94]. Development [WMTB94]. Device [TSSV94]. Directed [PM94]. Distributed [CK94]. Distributed-Memory [SSSK94]. Distribution [APG94].

Distribution-Independent [APG94]. Do [Gin94]. DOACROSS [CTY94]. DSM [MHS94]. Dynamic [KN94]. Dynamics [GC94].

Ecosystem [TT94].

Edinburgh [WMTB94]. Editor [HKTW94]. Education [MM94]. Efficient [KMY94]. Element [SHM^{+}94]. EM-4 [SSSK94]. Enable [TJH^{+}94]. Enabling [SM94]. Environment [BTS94]. Environmental [McR94]. Environments [Rov94]. Equation [NIM94]. Equations [BJTW94]. Evaluation [BY94]. Expectations [CPW94]. Experiences [CCFH94]. Experimental [BGKB94]. Exploratorium [IBRT94]. Expressing [ZI94]. expressions [BE94]. EXTENT [DGKL94].

Factorization [RS94]. Factors [WCL94]. Failures [GMB94]. Faster [McR94]. Fault [BC94]. Fault-Tolerant [BC94]. FFT [AGZ94]. Fibre [LHDM94]. File [KN94]. File-Access [KN94]. Finite [SHM^{+}94]. Fixed [BY94]. Flow [BTS94]. Flows [BJTW94]. Fluid [BTS94]. Formulation [KGK94]. Formulations [GKS94]. Frameworks [Rov94].

Garbage [KMY94]. Gauss [KRF94].

Generalized [PM94]. Generation [CCHZ94]. Generator [PCMR94]. Genetic [MMMR94]. Gflop [SHM^{+}94]. Gflop/s [SHM^{+}94]. Global [KMY94]. Gradient [TO94]. GRAPE [MTES94]. GRAPE-4 [MTES94]. Graph [MMMR94]. Graphs [PM94]. Grid [TSSV94]. Grids [BTS94].

Hardware [McR94]. High [Ben94, Lin94].

High-Performance [BTS94]. High-Speed [LHDM94]. HPCC [Lin94]. HPF [PHHF94]. HPF/Fortran [PHHF94]. Hut [GKS94]. Hypercube [KMB94]. Hyperplane [ZI94].

I/O [LHDM94]. IBM [Mra94]. Icosahedral [JWT^{+}94]. Image [TJH^{+}94]. Imaging [Fah94]. Implementation [PCB94]. Implementing [DGKL94]. Improved [CY94]. Improving [MHS94]. Incremental [MMMR94]. Independent [APG94]. Inference [KS94]. Information [Lin94]. Infrastructure [Lin94]. Initiative [HPS94]. Integer [Eck94]. Integrating [Rov94]. Integration [BGKB94]. Intel [WR94]. intensive [PP94]. Interactive [HKTW94]. Interface [PCB94]. Interior [KGK94]. Interpreting [PHHF94]. Intertask [CY94]. Involvement [MHS94]. IofNEWT [BGKB94]. iPSC [WR94]. iPSC/860 [WR94]. Irregular [TSSV94]. Issues [Ben94]. Iterative [TSSV94].

Joint [HPS94].

Latency [KBHA94]. Lauritzen [KS94]. Lauritzen-Spiegelhalter [KS94]. Level [FLRS94]. Library [PCB94]. Library-Based [PCB94]. Linear [TSSV94]. load [CKO^{+}94]. Local [LHDT94]. Locality [CY94]. Loop [ZI94]. Loops [Chi94]. Lost [CL94]. Low [KBHA94].

Machine [CK94, CWM94]. Machines [TO94]. Mammalian [WCL94]. Mapping [SOGD94]. Massively [KMY94]. Matrices [TSSV94]. Medicine [JP94]. Memory [CK94]. Mesh [BC94]. Meshes [CCHZ94]. Method [TO94]. Methods [TT94]. migration [CKO^{+}94].

Mixed [Eck94]. Model [NHL94]. Modeling [FR94]. Models [WCL94]. Molecular [GC94]. Monitoring [TT94]. Moving [CWM94]. MPP [Gin94]. Multicast [HKM94]. Multicomputers [LLWN94]. Multidimensional [SW94]. Multigrid [TO94]. Multiprocessor [PM94]. Multiprocessors [KMB94]. Multistage [TZ94].

NAL [MFI^{+}94]. NAS [CCFH94]. NASA [HPS94]. National [Lin94]. Navier [BJTW94]. Need [McR94]. Network [LHDT94, TZ94]. Networks [LHDT94]. Next [CCHZ94]. Node [GMB94]. Non [LLWN94]. Non-Adaptive [BC94]. Non-contiguous [LLWN94]. non-linear [BE94]. Nonnumeric [SSSK94]. November [IEE94]. NSF [HPS94]. NSF-NASA [HPS94]. Number [PCMR94]. Numerical [BGKB94]. NWT [MFI^{+}94].

O [LHDM94]. Object [PCB94]. Object-Oriented [PCB94]. Ocean [BJTW94]. One [MTES94]. One-Tflops [MTES94]. Only [GMB94]. Operation [WMTB94]. Operations [HKM94]. Optimal [XGN94]. Orientation [JWT^{+}94]. Oriented [PCB94]. OSF [SS94]. OSF/1 [SS94]. Other [McR94].

Pacemaker [WCL94]. Package [HCM94]. Paging [BHMW94]. Paragon [SS94, SHM^{+}94]. Parallel [BY94]. Parallelism [CPW94]. Parallelization [CTY94]. PARAMICS [CWM94]. Partitioning [MMMR94]. 860 [WR94]. Fortran [PHHF94]. or [TNC94]. S-15 [SS94]. Performance [Ben94].

Pipeline [KH94]. Planning [FR94]. Point [KGK94]. Portable [DGKL94]. Potential [CPW94]. Potentials [WCL94]. Power [KRF94]. Preconditioned [TO94]. Prediction [CL94]. Prefetch [Chi94]. Probabilistic [KS94]. Problem [APG94]. Problems [SPMH94]. Proceedings [IEE94]. Processor [LLWN94]. Production [KN94]. Program [Lin94]. Programmable [Chi94]. Programme [WMTB94]. Programming [DGKL94]. Programs [SOGD94]. Propagation [WCL94]. Protein [CSPA94]. Protocols [FLRS94]. Prototype [CCFH94]. Pseudorandom [PCMR94]. Purpose [MTES94]. PVM [CKO^{+}94].

Range [BE94]. Reacting [SHM^{+}94]. Real [Ben94]. Real-Time [Ben94]. Realizing [McR94]. Really [Gin94]. Reconstruction [JWT^{+}94]. Recursive [DGKL94]. Reducing [Mra94]. Reproducible [PCMR94]. Requirements [SOGD94]. Research [FR94]. Results [SSSK94]. Routed [XGN94]. Routing [BC94]. Run [SPMH94]. Run-time [SPMH94].

s [SHM^{+}94]. SA [WCL94]. Scalability [TV94]. Scalable [BTS94]. Scheduling [SE94]. Scheme [CY94]. Schemes [KMB94]. Scholarship [WMTB94]. Science [MM94]. Scientific [FW94]. Search [SSSK94]. Seidel [KRF94]. Selecting [Gin94]. Self [POVDF94]. Self-Compacting [POVDF94]. Semiconductor [TSSV94]. Server [TJH^{+}94]. Shared [NHL94]. Shared-Memory [NHL94]. Sharing [TNC94]. Simulation [SHM^{+}94]. Simulations [GKS94]. SLICC [KBHA94]. Software [XGN94]. Solver [TSSV94].

Solving [TSSV94]. Space [TNC94]. Sparse [KRF94]. Special [MTES94]. Special-Purpose [MTES94]. Specific [FLRS94]. Speed [LHDM94]. Spiegelhalter [KS94]. Steering [JP94]. Stokes [BJTW94]. Story [Gin94]. Strategies [Eck94]. Structure [CSPA94]. Study [LHDM94]. Subsystems [LHDM94]. Summer [WMTB94]. Sunder [Chi94]. SUNMOS [SS94]. Supercomputer [UIT94]. Supercomputers [KSF^{+}94]. Supercomputing [Fah94, IEE94]. Support [SPMH94]. Survey [CPW94]. Switching [TZ94]. symbolic [BE94]. System [KRF94]. Systems [Ben94].

Talking [HCM94]. Task [SOGD94].

Techniques [IBRT94].

Technologies [Lin94]. Tell [Gin94]. test [BE94]. Tflops [MTES94]. Threads [HCM94]. Three [FW94]. Tight [GC94]. Time [Ben94]. Tolerant [BC94]. Tolerating [GMB94]. Too [CPW94]. Tool [HKTW94]. Tools [Rov94]. Tradeoffs [BHMW94]. Transfers [Mra94]. Transformations [SM94]. Truly [APG94]. Tunnel [MFI^{+}94]. Tunnels [BGKB94].

Unbalanced [SE94]. Uncertain [CSPA94]. Undergraduate [MM94]. Unimodular [SM94]. Unstructured [BTS94]. User [FLRS94]. User-Level [FLRS94]. Users [CPW94]. Using [BJTW94].

Variance [Mra94]. Vector [KSF^{+}94]. Vehicles [CWM94]. via [MHS94]. Viruses [JWT^{+}94]. VPP500 [UIT94].

Washington [IEE94]. Whole [Gin94]. Wind [BGKB94]. Workload [KN94]. Workloads [SE94]. Workstation [TNC94].

Workstations [CCFH94]. Wormhole [BC94]. Wormhole-Routed [XGN94].

XP [SS94]. XP/S [SS94]. XP/S-15 [SS94].

References

Agarwal:1994:HPP

[AGZ94] R. C. Agarwal, F. G. Gustavson, and M. Zubair. A high performance parallel algorithm for 1-D FFT. In IEEE [IEE94], pages 34–40. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Aluru:1994:TDI

[APG94] S. Aluru, G. M. Prabhu, and J. Gustafson. Truly distributionindependent algorithms for the N-body problem. In IEEE [IEE94], pages 420–428. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Boppana:1994:FTR

[BC94] Rajendra V. Boppana and Suresh Chalasani. Fault-tolerant routing with non-adaptive wormhole algorithms in mesh networks. In IEEE [IEE94], pages 693–702. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Blume:1994:RTD

[BE94] William Blume and Rudolf Eigenmann. The range test: a dependence test for symbolic, non-linear expressions. In IEEE [IEE94], pages 528–537. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Benincasa:1994:IAH

[Ben94] M. Benincasa. Issues in applying high performance computing to real-time systems. In IEEE [IEE94], pages 1–2. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Bailey:1994:INE

[BGKB94] F. R. Bailey, M. George, D. Koga, and P. Buning. Integration of Numerical and Experimental Wind Tunnels (IofNEWT). In IEEE [IEE94], page 7. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.

gov/AP/contents.html. IEEE

catalog number 94CH34819.

Barnett:1994:BHP

[BGPS94] M. Barnett, S. Gupta, D. G.

Payne, and L. Shuler. Building a high-performance collective communication library. In IEEE [IEE94], pages 107–116. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Burger:1994:PTD

[BHMW94] Douglas C. Burger, Rahmat S.

Hyder, Barton P. Miller, and David A. Wood. Paging tradeoffs in distributed sharedmemory multiprocessors. In IEEE [IEE94], pages 590–599. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Beddhu:1994:TCO

[BJTW94] M. Beddhu, M.-Y. Jiang, L. Taylor, and D. Whitfield. Towards computations of ocean flows using Navier–Stokes equations. In IEEE [IEE94], pages 144–153. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/

contents.html. IEEE catalog

number 94CH34819.

Banerjee:1994:SHP

[BTS94] D. Banerjee, T. Tysinger, and W. Smith. A scalable highperformance environment for fluid flow analysis on unstructured grids. In IEEE [IEE94], pages 8–17. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Baker:1994:FTP

[BY94] A. M. Baker and S. X. Ying. A fixed time performance evaluation of parallel CFD applications. In IEEE [IEE94], pages 18–23. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Castagnera:1994:NEP

[CCFH94] K. Castagnera, D. Cheng, R. Fatoohi, and E. Hook. NAS experiences with a. prototype cluster of workstations. In IEEE [IEE94], pages 410–419. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Christon:1994:MNG

[CCHZ94] M. Christon, J. Compton, D. Hardin, and M. Zosel. Meshes: The next generation. In IEEE [IEE94], pages 275–276. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Cheng:1994:PDP

[CH94] Doreen Cheng and Robert Hood.

A portable debugger for parallel and distributed programs. In IEEE [IEE94], pages 723–732. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Chiueh:1994:SPH

[Chi94] T.-C. Chiueh. Sunder: A programmable hardware prefetch architecture for numerical loops. In IEEE [IEE94], pages 488–497. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Ching:1994:EAC

[CK94] W.-M. Ching and A. Katz. An experimental APL compiler for a distributed memory parallel machine. In IEEE [IEE94], pages 59–68. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Casas:1994:ALM

[CKO^{+}94] J. Casas, R. Konuru, S. W. Otto, R. Prouty, and J. Walpole. Adaptive load migration systems for PVM. In IEEE [IEE94], pages 390–399. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Crovella:1994:PPP

[CL94] Mark E. Crovella and Thomas J. LeBlanc. Parallel performance prediction using lost cycles analysis. In IEEE [IEE94], pages 600– 609. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Cook:1994:EPT

[CPW94] C. R. Cook, C. M. Pancake, and R. Walpole. Are expectations for parallelism too high? A survey of potential parallel users. In IEEE [IEE94], pages 126–133. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Chen:1994:PPS

[CSPA94] Cheng Che Chen, Jaswinder Pal Singh, William B. Poland, and Russ B. Altman. Parallel protein structure determination from uncertain data. In IEEE [IEE94], pages 570–579. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Chen:1994:EAR

[CTY94] Ding-Kai Chen, Josep Torrellas, and Pen-Chung Yew. An efficient algorithm for the run-time parallelization of DOACROSS loops. In IEEE [IEE94], pages 518–527. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Cameron:1994:PMV

[CWM94] G. Cameron, B. J. N. Wylie, and D. McArthur. PARAM-ICS: Moving vehicles on the Connection Machine. In IEEE [IEE94], pages 291–300. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Choi:1994:CDC

[CY94] Lynn Choi and Pen-Chung Yew. A compiler-directed cache coherence scheme with improved intertask locality. In IEEE [IEE94], pages 773–782. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Dai:1994:EPP

[DGKL94] D. L. Dai, S. K. S. Gupta, S. D. Kaushik, and J. H. Lu. EX-TENT: A portable programming environment for designing and implementing high-performance block-recursive algorithms. In IEEE [IEE94], pages 49–58. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Eckstein:1994:CSP

[Eck94] J. Eckstein. Control strategies for parallel mixed integer branch and bound. In IEEE [IEE94], pages 41–48. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535.

LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Fahlman:1994:SSA

[Fah94] S. Fahlman. Supercomputing support for advanced biomedical imaging. In IEEE [IEE94], pages 538–539. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Falsafi:1994:ASP

[FLRS94] B. Falsafi, A. R. Lebeck, S. K.

Reinhardt, and I. Schoinas. Application-specific protocols for user-level shared memory. In IEEE [IEE94], pages 380–389. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Feurzeig:1994:CMS

[FR94] W. Feurzeig and N. Roberts.

Computer modeling in science education: Toward a research and planning agenda. In IEEE [IEE94], pages 5–6. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/

                                        contents.html. IEEE catalog

number 94CH34819.

Fatoohi:1994:PET

[FW94] R. Fatoohi and S. Weeratunga. Performance evaluation of three distributed computing environments for scientific applications. In IEEE [IEE94], pages 400–409. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Goedecker:1994:TBM

[GC94] S. Goedecker and L. Colombo. Tight binding molecular dynamics. In IEEE [IEE94], pages 670– 672. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

                                        URL http://sc94.ameslab.
                                        gov/AP/contents.html. IEEE

catalog number 94CH34819.

Ginsberg:1994:SBM

[Gin94] M. Ginsberg. Selecting the “best” MPP: Do benchmarks really tell the whole story? In IEEE [IEE94], page 4. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Gupta:1994:SPA

[GK94] Anshul Gupta and Vipin Kumar.

A scalable parallel algorithm for sparse Cholesky factorization. In IEEE [IEE94], pages 793–802. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Grama:1994:SPF

[GKS94] A. Y. Grama, V. Kumar, and A. Sameh. Scalable parallel formulations of the Barnes–Hut method for n-body simulations. In IEEE [IEE94], pages 439–448. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Gefflaut:1994:TNF

[GMB94] A. Gefflaut, C. Morin, and M. Banatre. Tolerating node failures in cache only memory architectures. In IEEE [IEE94], pages 370–379. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Haines:1994:DCT

[HCM94] M. Haines, D. Cronk, and P. Mehrotra. On the design of Chant: A talking threads package. In IEEE [IEE94], pages 350– 359. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Huang:1994:DIM

[HKM94] C. Huang, E. P. Kasten, and P. K. McKinley. Design and implementation of multicast operations for ATM-Based high performance computing. In IEEE [IEE94], pages 164–173. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Hiranandani:1994:DEN

[HKTW94] Seema Hiranandani, Ken Kennedy, Chau-Wen Tseng, and Scott Warren. The D editor: A new interactive parallel programming tool. In IEEE [IEE94], pages 733–742. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Hotovy:1994:JNN

[HPS94] S. Hotovy, W. Pfeiffer, and T. Sterling. Joint NSF-NASA initiative in evaluation. In IEEE [IEE94], pages 811–813. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Iyer:1994:DEH

[IBRT94] S. V. Iyer, A. Beck, U. Ravaioli, and J. Terstriep. Distributed exploratorium for high performance computational techniques. In IEEE [IEE94], pages 117–125. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

IEEE:1994:PSW

[IEE94] IEEE, editor. Proceedings, Supercomputing ’94: Washington, DC, November 14–18, 1994, Supercomputing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Johnson:1994:CSM

[JP94] Christopher R. Johnson and Steven G. Parker. A computational steering model applied to problems in medicine. In IEEE [IEE94], pages 540–549. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Johnson:1994:ODR

[JWT^{+}94] C. A. Johnson, N. I. Weisenfeld, B. L. Trus, J. F. Conway, R. L. Martino, and A. C. Steven. Orientation determination in the 3D reconstruction of icosahedral viruses using a parallel computer. In IEEE [IEE94], pages 550–559. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Knies:1994:SLL

[KBHA94] A. D. Knies, F. R. Barriuso, W. J. Harrod, and G. B. Adams. SLICC: A low latency interface for collective communications. In IEEE [IEE94], pages 89–96. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Karypis:1994:PFI

[KGK94] G. Karypis, A. Gupta, and V. Kumar. A parallel formulation of interior point algorithms.

In IEEE [IEE94], pages 204–213. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Kee:1994:ECA

[KH94] K. K. Kee and S. Hariri. Efficient communication algorithms for pipeline multicomputers. In IEEE [IEE94], pages 468–477. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Kumar:1994:ESC

[KMB94] A. Kumar, P. K. Mannava, and L. N. Bhuyan. Efficient and scalable cache coherence schemes for shared memory hypercube multiprocessors. In IEEE [IEE94], pages 498–507. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Kamada:1994:EPG

[KMY94] T. Kamada, S. Matsuoka, and A. Yonezawa. Efficient parallel global garbage collection on massively parallel computers. In IEEE [IEE94], pages 79–88. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Kotz:1994:DFA

[KN94] David Kotz and Nils Nieuwejaar. Dynamic file-access characteristics of a production parallel scientific workload. In IEEE [IEE94], pages 640–649. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Koester:1994:PGS

[KRF94] D. P. Koester, S. Ranka, and G. C. Fox. A parallel Gauss– Seidel algorithm for sparse power system matrices. In IEEE [IEE94], pages 184–193. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Kozlov:1994:PLS

[KS94] A. V. Kozlov and Jaswinder Pal Singh. A parallel Lauritzen-Spiegelhalter algorithm for probabilistic inference. In IEEE [IEE94], pages 320–329. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535.

LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Kontothanassis:1994:CPV

[KSF^{+}94] L. Kontothanassis, R. A. Sugumar, G. Faanes, J., and J. E. Smith. Cache performance in vector supercomputers. In IEEE [IEE94], pages 255–264. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Lin:1994:PHS

[LHDM94] M. Lin, J. Hsieh, D. H. C. Du, and J. A. MacDonald. Performance of high-speed network I/O subsystems for: Case study of a fibre channel network. In IEEE [IEE94], pages 174–183. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Lin:1994:DNC

[LHDT94] M. Lin, J. Hsieh, D. H. C.

Du, and J. P. Thomas. Distributed network computing over local ATM networks. In IEEE [IEE94], pages 154–163. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Lindberg:1994:HPC

[Lin94] D. Lindberg. The High Performance Computing and Communications (HPCC) program: Technologies for the National Information Infrastructure. In IEEE [IEE94], page 279. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Liu:1994:NCP

[LLWN94] W. Liu, V. Lo, K. Windisch, and B. Nitzberg. Non-contiguous processor allocation algorithms for distributed memory multicomputers. In IEEE [IEE94], pages 227–236. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

McRae:1994:RBH

[McR94] G. J. McRae. Realizing the benefits of high performance computing for solving environmental and other problems: What do we need beyond faster hardware? In IEEE [IEE94], page 280. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535.

LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Miyoshi:1994:DAN

[MFI^{+}94] H. Miyoshi, M. Fukuda, T. Iwamiya, T. Nakamura, M. Tuchiya, M. Yoshida, K. Yamamoto, Y. Yamamoto, S. Ogawa, Y. Matsuo, T. Yamane, M. Takamura, M. Ikeda, S. Okada, Y. Sakamoto, et al. Development and achievement of NAL Numerical Wind Tunnel (NWT) for CFD computations. In IEEE [IEE94], pages 685–692. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Mirchandaney:1994:IPD

[MHS94] Ravi Mirchandaney, Seema Hiranandani, and Ajay Sethi. Improving the performance of DSM systems via compiler involvement. In IEEE [IEE94], pages 763–772. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Marchioro:1994:IUE

[MM94] T. Marchioro and D. Martin.

Issues in undergraduate education in computational science and high performance computing. In IEEE [IEE94], page 3. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Maini:1994:GAG

[MMMR94] H. S. Maini, K. G. Mehrotra, C. K. Mohan, and S. Ranka. Genetic algorithms for graph partitioning and incremental graph partitioning. In IEEE [IEE94], pages 449–457. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Mraz:1994:RVP

[Mra94] Ronald Mraz. Reducing the variance of point to point transfers in the IBM 9076 parallel computer. In IEEE [IEE94], pages 620–629. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Makino:1994:GOT

[MTES94] J. Makino, M. Taiji, T. Ebisuzaki, and D. Sugimoto. GRAPE-4: A one-tflops special-purpose computer for astrophysical Nbody problem. In IEEE [IEE94], pages 429–438. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Nieplocha:1994:GAP

[NHL94] J. Nieplocha, R. J. Harrison, and R. J. Littlefield. Global arrays: A portable “shared-memory” programming model for distributed memory computers. In IEEE [IEE94], pages 340–349. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Nakanishi:1994:HPL

[NIM94] Makoto Nakanishi, Hiroshi Ina, and Kenichi Miura. A high performance linear equation solver on the VPP500 parallel supercomputer. In IEEE [IEE94], pages 803–810. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Ou:1994:PIG

[OR94] C.-W. Ou and S. Ranka. Parallel incremental graph partitioning using linear programming. In IEEE [IEE94], pages 458–467.

ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Parkes:1994:LBA

[PCB94] S. Parkes, J. A. Chandy, and P. Banerjee. A library-based approach to portable, parallel, object-oriented programming: Interface, implementation, and application. In IEEE [IEE94], pages 69–78. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Pryor:1994:IPR

[PCMR94] D. V. Pryor, S. A. Cuccaro, M. Mascagni, and M. L. Robinson. Implementation of a portable and reproducible parallel pseudorandom number generator. In IEEE [IEE94], pages 311–319. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Parashar:1994:IPH

[PHHF94] Manish Parashar, Salim Hariri, Tomasz Haupt, and Geoffrey C. Fox. Interpreting the performance of HPF/Fortran 90D. In IEEE [IEE94], pages 743–752. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Prasanna:1994:GMS

[PM94] G. N. Srinivasa Prasanna and B. R. Musicus. Generalized multiprocessor scheduling for directed acyclic graphs. In IEEE [IEE94], pages 237–246. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Park:1994:DED

[POVDF94] J. Park, B. W. O’Krafka, S. Vassiliadis, and J. Delgado-Frias. Design and evaluation of a DAMQ multiprocessor network with self-compacting buffers. In IEEE [IEE94], pages 713–722. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Pasquale:1994:DCI

[PP94] Barbara K. Pasquale and George C.

Polyzos. Dynamic I/O characterization of I/O intensive scientific applications. In IEEE [IEE94], pages 660–669. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Rover:1994:PEI

[Rov94] D. Rover. Performance evaluation: Integrating techniques and tools into environments and frameworks. In IEEE [IEE94], pages 277–278. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Rothberg:1994:ILD

[RS94] Edward Rothberg and Robert Schreiber. Improved load distribution in parallel sparse Cholesky factorization. In IEEE [IEE94], pages 783–792. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Subramaniam:1994:ASU

[SE94] S. Subramaniam and D. Eager. Affinity scheduling of unbalanced workloads. In IEEE [IEE94], pages 214–226. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Shadid:1994:GUF

[SHM^{+}94] John N. Shadid, Scott A.

Hutchinson, Harry K. Moffat, Gary L. Hennigan, Bruce Hendrickson, and Robert W. Leland. A 65+ Gflop/s unstructured finite element simulation of chemically reacting flows on the Intel Paragon. In IEEE [IEE94], pages 673–679 (or 673–678??). ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Sass:1994:EUT

[SM94] Ron Sass and Matt Mutka.

Enabling unimodular transformations. In IEEE [IEE94], pages 753–762. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Subhlok:1994:CMR

[SOGD94] J. Subhlok, D. R. O’Hallaron, T. Gross, and P. A. Dinda. Communication and memory requirements as the basis for mapping task and data parallel programs. In IEEE [IEE94], pages 330–339.

ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Sharma:1994:RTC

[SPMH94] S. D. Sharma, R. Ponnusamy, B. Moon, and Y.-S. Hwang. Runtime and compile-time support for adaptive irregular problems. In IEEE [IEE94], pages 97–106. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Saini:1994:APU

[SS94] Subhash Saini and Horst D. Simon. Applications performance under OSF/1 AD and SUNMOS on Intel Paragon XP/S-15. In IEEE [IEE94], pages 580–589. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Sato:1994:NSR

[SSSK94] M. Sato, A. Sohn, S. Sakai, and Y. Kodawa. Nonnumeric search results on the EM-4 distributedmemory multiprocessor. In IEEE [IEE94], pages 301–310. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Seamons:1994:EAI

[SW94] Kent E. Seamons and Marianne Winslett. An efficient abstract interface for multidimensional array I/O. In IEEE [IEE94], pages 650–659. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Tierney:1994:UHS

[TJH^{+}94] Brian L. Tierney, William E.

Johnston, Hanan Herzog, Gary Hoo, Guojun Jin, Jason Lee, Ling Tony Chen, and Doron Rotem. Using high speed networks to enable distributed parallel image server systems. In IEEE [IEE94], pages 610–619. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Turner:1994:TAS

[TNC94] Stephen W. Turner, Lionel M.

Ni, and Betty H. C. Cheng. Time and/or space sharing in a workstation cluster environment. In IEEE [IEE94], pages 630–639. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Tatebe:1994:EIM

[TO94] O. Tatebe and Y. Oyanagi. Efficient implementation of the multigrid preconditioned conjugate gradient method on distributed memory machines. In IEEE [IEE94], pages 194–203. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Tomacruz:1994:PIL

[TSSV94] E. Tomacruz, J. Sanghavi, Sangiovanni, and A. Vincentelli. A parallel iterative linear solver for solving irregular grid semiconductor device matrices. In IEEE [IEE94], pages 24–33. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.

Turner:1994:ADP

[TT94] C. J. Turner and J. G. Turner. Adaptive data parallel methods for ecosystem monitoring. In IEEE [IEE94], pages 281–290. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Turner:1994:SCS

[TV94] S. W. Turner and A. V. Viedenbaum. Scalability of the Cedar system. In IEEE [IEE94], pages 247–254. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Torrellas:1994:PCM

[TZ94] J. Torrellas and Z. Zhang. The performance of the Cedar Multistage Switching Network. In IEEE [IEE94], pages 265–274. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Utsumi:1994:AVP

[UIT94] T. Utsumi, M. Ikeda, and M. Takamura. Architecture of the VPP500 parallel supercomputer. In IEEE [IEE94], pages 478–487. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/

contents.html. IEEE catalog

number 94CH34819.

Winslow:1994:FCG

[WCL94] Raimond L. Winslow, Dongming Cai, and Ying-Cheng Lai. Factors controlling generation and propagation of pacemaker potentials in network models of mammalian SA node. In IEEE [IEE94], pages 560–569. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Womble:1994:ABE

[WGW^{+}94] David E. Womble, David S. Greenberg, Stephen R. Wheat, Robert E. Benner, Marc S. Ingber, Greg Henry, and Satya Gutpa. Applications of boundary element methods on the Intel Paragon. In IEEE [IEE94], pages 680–684. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Wilson:1994:DOE

[WMTB94] G. V. Wilson, N. B. MacDonald, C. Thornborrow, and C. M. Brough. The development and operation of Edinburgh Parallel Computing Centre’s summer scholarship programme. In IEEE [IEE94], pages 134–143.

ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Wang:1994:SUC

[WR94] J.-C. Wang and S. Ranka.

Scheduling of unstructured communication on the Intel iPSC/ 860. In IEEE [IEE94], pages 360– 369. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Xu:1994:OSM

[XGN94] Hong Xu, Ya-Dong Gui, and Lionel M. Ni. Optimal software multicast in wormholerouted multistage networks. In IEEE [IEE94], pages 703–712. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL

http://sc94.ameslab.gov/AP/
contents.html. IEEE catalog

number 94CH34819.

Zaafrani:1994:ECL

[ZI94] A. Zaafrani and M. R. Ito. Expressing cross-loop dependencies through hyperplane data dependence analysis. In IEEE [IEE94], pages 508–517. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535.

LCCN QA76.5 .S894 1994. URL

                                        http://sc94.ameslab.gov/AP/
                                        contents.html. IEEE catalog

number 94CH34819.